# RAGONITE: Iterative Retrieval on Induced Databases and Verbalized RDF for Conversational QA over KGs with RAG

### v1

@@
Abstract questions like ‘Innovative highlights in x7?’ require NL understanding and common sense reasoning to map the intent to equipments and accessories of the BMW X7. We address such questions by verbalizing KGs [Og22], so that LLMs can reason over them like NL text. To do this, we use the KG in Turtle format, that encapsulates all facts of a specific subject entity. The facts in each “capsule” are then verbalized into NL passages using a simple rule-based system: (i) all schema-related prefixes are stripped to extract the raw content; (ii) the subject string, predicate string and object string are concatenated to form a sentence; (iii) “is” is added before type predicates, and “has” before other predicates; and (iv) The reverse formulation of the fact is also added to the passage to later help the generator LLM answer questions when the contained information is requested in complementary ways (What is BMW 120 Sport’s engine performance? and Which BMW engines have a performance of 125 kW?). Explicitly adding reversed facts is helpful for smaller LLMs.

### v2

@@
Abstract questions like ‘Innovative highlights in x7?’ require NL understanding and common sense reasoning to map the intent to equipments and accessories of the BMW X7. We address such questions by verbalizing KGs [Og22], so that LLMs can reason over them like NL text. We again use the KG in Turtle format, where all facts with a specific entity as subject are grouped together. The facts in each Turtle “capsule” are then verbalized into NL passages using simple rules: (i) all schema-related prefixes are stripped to extract the raw content; (ii) the subject string, predicate string and object string are concatenated to form a sentence; (iii) “is” is added before type predicates, and “has” before other predicates; and (iv) The reverse formulation of the fact is also added to the passage to later help the generator LLM answer questions when the contained information is requested in complementary ways (What is BMW 120 Sport’s engine performance? and Which BMW engines have a performance of 125 kW?). Explicitly adding reversed facts is helpful for smaller LLMs.